i have a math problem that i dont get that says the perimeter of a rectangle is 126 centimeters.the lenght is nine centimeters longer than the width.what are the dimensions of the rectangle
Answer by stanbon(75887)

the perimeter of a rectangle is 126 centimeters.
2(length + width) = 126
----------------------
the length is nine centimeters longer than the width.
Let the width be "W"
Then the length is "W+9"
-----
what are the dimensions of the rectangle
-----
Sustitute for the length and wide and solve:
2(W+9 + W) = 126
Divide by 2
2W + 9 = 63
Subtract 9
2W = 54
Divide by 2
W = 27 cm (the width)
---
Solve for length:
W+9 = 36 cm (the length)
============================
Cheers,
Stan H.
==========